# Binary population synthesis of the Galactic canonical pulsar population

### v1

@@ 2.3 COMPAS Implementation of Pulsar Evolution
This study focuses on canonical pulsars, hence we only describe the evolution of isolated pulsars. Here we follow magnetic braking for an isolated pulsar to spin down, with braking index $n=3$. The pulsar evolution theory implemented in COMPAS is described in Chattopadhyay et al. (2020). We assume an exponentially decaying magnetic field of a neutron star after it spins down for a time duration of $t$
@@
where $B_{i}$ is the magnetic field strength of the neutron star before it is evolved by time duration $t$ and $B_{f}$ the magnetic field strength afterwards, $B_{\textrm{min}}$ is the minimum magnetic field of a pulsar which is set to be $10^{8}$ G for the entire analysis, and $\tau$ is the time scale of pulsar magnetic field decay.
During the time duration $t$, the change in the spin period of the neutron star is described as
$P^{2}=\frac{16\pi^{2}R^{6}\sin^{2}{\alpha}}{3c^{3}I}[B_{\textrm{min}}^{2}t-\tau B_{\textrm{min}}(B_{f}-B_{i})-\frac{\tau}{2}(B^{2}_{f}-B^{2}_{i})]+P_{0}^{2},$ (3)
where $R$ is the radius of the pulsar (we assume $R=10$ km for all pulsars at all times in this work), $P_{0}$ is the spin period of the pulsar at the beginning of the evolved time duration $t$, and $B_{i}$ and $B_{f}$ are the pulsar surface magnetic field at the beginning and end of the time step as described in Eq. 2, $c$ is speed of light, $\alpha$ is the alignment angle between the magnetic and rotational axes, and
$I=(0.237\pm 0.008)mR^{2}[1+4.2\frac{\textrm{M}}{\textrm{M}_{\odot}}\frac{\textrm{km}}{\textrm{R}}+90(\frac{\textrm{M}}{\textrm{M}_{\odot}}\frac{\textrm{km}}{\textrm{R}})^{4}]$ (4)
being the moment of inertia of the pulsar (Lattimer & Schutz, 2005). Currently in COMPAS, $\alpha=\pi/2$ is set to be constant and not evolved (though see Johnston et al., 2020). Evolution of alignment will be included in a future version of COMPAS.
The evolution of period derivative, $\dot{P}$ , can be described after updating magnetic field strength and spin period, as
@@ 2.4 Free Parameters of COMPAS Runs
In addition to the default binary parameters, we need to propose distributions of pulsar properties at birth for subsequent modelling. We note that some of the past population studies (e.g., Faucher-Giguère & Kaspi, 2006; Popov & Turolla, 2012; Szary et al., 2014; Johnston et al., 2020; Chattopadhyay et al., 2020; Dirson et al., 2022; Igoshev et al., 2022) proposed and refined distributions of pulsar properties at birth, such as surface magnetic field strength and spin period. In this work, we also choose the distributions of these two quantities at birth, as well as magnetic decay timescale ($\tau_{d}$) to be free parameters. By default, COMPAS sets the birth distribution of NS magnetic field as a uniform distribution between $10^{10}$ G to $10^{13}$ G, and the birth distribution of NS spin period as a uniform distribution between 10 to 100 ms. These distributions are in general agreement with those observed from young pulsar populations (Manchester et al., 2005). We exclude extremely high magnetic field strengths with B $>10^{15}$ G in the magnetar regime as we are uncertain if the magnetar formation and evolution physics would follow those of a normal NS. In our default model, we assume that the magnetic field decay time scale $\tau_{d}=500$ Myr (Chattopadhyay et al., 2021). We choose $\tau_{d}$ as a free parameter since there is no clear consensus on whether the magnetic field of a pulsar decays at all, or how long it takes to decay. For example, semi-analytical studies (Gunn & Ostriker, 1970; Vivekanand & Narayan, 1981; Lyne et al., 1985) found that the magnetic field decays on the timescale of the order of a million years, with Dirson et al. (2022) recently finding $\tau_{d}=4.6\times 10^{5}$ yr. In comparison, other works (Taylor & Manchester, 1977; Stollman, 1987; Kiel et al., 2008) have shown the opposite, finding that $\tau_{d}>100$ Myr is needed in order for theoretical models to match the observed catalog. For this reason, we let $\tau_{d}$ vary from 10 Myr to 1000 Myr in our models.
@@
We list the details of all the models with different pulsar birth distributions and parameters in Table. 1. The initial model (INIT) is as described above. Following that, we vary our model parameters in such ways that we cover different ranges, and in some cases different shapes of the distribution. The D1, D2 and D3 models, combined with INIT, are specifically targeted to examine the impact of the $\tau_{d}$ on the results of the simulations, as only $\tau_{d}$ is varied in models D1, D2 and D3 compared to INIT. For B1 and B2 models, only the range of the birth magnetic field distributions is changed compared to INIT. Similarly for P1, P2 and P3 models, only the ranges of the birth spin period distributions are changed and the others are kept the same. For PN1, PN2 and PN3 models, the birth spin period distribution is changed to a normal distribution (hence, PN models as in period-normal models) with various means and standard deviations. In the BLN series of models, the birth magnetic field strength distribution is changed to a log normal distribution, with different means and standard deviations in models BLN1–4. The remaining three models, BLN1-PN1, BLN4-PN1, and BLN1-PN5, have a combination of the magnetic field distribution from the BL models and the spin period distributions of the PN models, respectively. We specifically note that the original form of the log-normal distribution of the birth magnetic field strength is taken from Szary et al. (2014). The BLN1-PN5 model is exactly as described in Szary et al. (2014). In each COMPAS model, 300,000 binary systems are modelled.
@@
Initial ($10^{10}$ $10^{13}$) Uniform (10 - 100) Uniform 500
D1 ($10^{10}$ $10^{13}$) Uniform (10 - 100) Uniform 100
D2 ($10^{10}$ $10^{13}$) Uniform (10 - 100) Uniform 10
D3 ($10^{10}$ $10^{13}$) Uniform (10 - 100) Uniform 1000
B1 ($10^{11}$ $10^{13}$) Uniform (10 - 100) Uniform 500
B2 ($10^{10}$ $10^{12}$) Uniform (10 - 100) Uniform 500
P1 ($10^{10}$ $10^{13}$) Uniform (10 - 1000) Uniform 500
P2 ($10^{10}$ $10^{13}$) Uniform (1 - 100) Uniform 500
P3 ($10^{10}$ $10^{13}$) Uniform (1 - 1000) Uniform 500
PN1 ($10^{10}$ $10^{13}$) Uniform $\mu=75$,$\sigma=25$ Normal 500
PN2 ($10^{10}$ $10^{13}$) Uniform $\mu=50$,$\sigma=25$ Normal 500
PN3 ($10^{10}$ $10^{13}$) Uniform $\mu=50$,$\sigma=50$ Normal 500
@@
PN1-B1 ($10^{11}$ $10^{13}$) Uniform $\mu=75$,$\sigma=25$ Normal 500
BLN1-PN1 $\mu=10^{12.6},\sigma=10^{0.55}$ LogNormal $\mu=75$,$\sigma=25$ Normal 500
BLN4-PN1 $\mu=10^{12.0},\sigma=10^{0.25}$ LogNormal $\mu=75$,$\sigma=25$ Normal 500
BLN1-PN5 $\mu=10^{12.6},\sigma=10^{0.55}$ LogNormal $\mu=300$,$\sigma=150$ Normal 500
@@ 2.5 Binary Birth, Pulsar Death, & Parameters
COMPAS begins evolving each binary system at $T=0$ Myr. The first stage of post-processing is to assign a birth time of the binary system that corresponds to the Milky Way star formation history. We consider a uniform star formation history of the Milky Way (Vigna-Gómez et al., 2018), enabling us to draw the birth time of the binary systems from a uniform distribution. Without recycling, pulsars rapidly spin down, losing their rotational energy to the point that they can no longer sustain detectable radio or $\gamma$-ray emission, at which point we define this as the death of pulsar. The typical characteristic age of a canonical $\gamma$-ray pulsar is relatively small, roughly $10^{5}$ yrs (Smith et al., 2023). For this reason, we assume that all the neutron stars from the simulation that would potentially be $\gamma$-ray emitting are formed in binary systems that are born in the recent history of the Milky Way. The birth time of any given binary system is then drawn from a uniform distribution with a minimum of 12.8 Gyr and a maximum of 13.0 Gyr. We choose to neglect the population of pulsars produced through accretion-induced collapse of white dwarfs as they represent only a small fraction of the population ($\lesssim 10\%$), are formed on much longer timescales and may potentially be born with different properties to those born in core-collapse supernovae (Bailyn & Grindlay, 1990; Tauris et al., 2013).
@@
Before recording the parameters of the neutron star, it is essential to check if the pulsar crosses the radio death-line (Rudak & Ritter, 1994) on the $P$-$\dot{P}$ diagram, which is defined by
@@
If $\dot{P}$ of a neutron star from COMPAS is larger than both of these two calculated values according to Eqs. 6,  7, then the pulsar has not yet crossed the death-line and may be observable, and we record its parameters. Our treatment of radio selection effects is given in Section 2.7.
@@ 2.6 Dynamical Evolution
Massive stars, which in later evolutionary stages form neutron stars, are formed within an exponential disk around the Galactic plane. We assume that these massive stars move along their Galactic orbits without any interruption until the change in the evolutionary status of one of the stars in the binary system. As described in § 2.1, we assume all the neutron stars are formed in supernovae. Upon formation, the stellar remnant receives a natal kick as described in § 2.2 which can disrupt the binary system, and in some cases, end up with two individual stars.
Upon the supernova event that forms the neutron star, the binary system is then assigned a random location in the Galaxy following a distribution of a thin, exponential disk (Paczynski, 1990):
$p_{z}(z)dz=\exp{-z/z_{\textrm{exp}}}\frac{dz}{z_{\textrm{exp}}},$ (8)
where $z$ is the vertical distance from the Galactic plane, and
$p_{R}(R)dR=a_{R}\exp{-R/R_{\textrm{exp}}}\frac{R}{R^{2}_{\textrm{exp}}}dR,$ (9)
where $a_{R}=1.0683$ and $R$ is the distance from the Galactic center on the Galactic plane. We set $z_{\textrm{exp}}=75$ pc and $R_{\textrm{exp}}=4.5$ kpc following Paczynski (1990).
Knowing the location of the newly born neutron star, the time at which it is born, and both the magnitude and direction of the natal kick velocity, we can follow its subsequent motion in the Galactic potential. We use NIGO (Rossi, 2015), a numerical galactic orbit integration tool, to evolve the motion of binary systems after the supernova within the Galactic potential. NIGO describes the Milky Way as a disc galaxy, with the Galactic potential comprised of three components: a bulge, a disk and a dark matter halo. The Galactic bulge is described as a Plummer sphere (Flynn et al., 1996; Plummer, 1911):
@@
where $G$ is the gravitation constant, $M_{b}=1.0\times 10^{10}$ is the mass of the bulge and $b_{b}=0.4$kpc is the scale length and $R^{2}=x^{2}+y^{2}$.
The Galactic disc is modelled by an exponential disc formed by the superposition of three Miyamoto-Nagai potentials (Miyamoto & Nagai, 1975):
@@
where $M_{d_{n}}$ are the masses of each disk, $a_{d_{n}}$ are related to the disk scale lengths of the three disk components (which are listed in Table. 2) and $b_{d}=0.2376$ is related to the disk scale height.
@@
$M_{d}$ (M) $9.0335\times 10^{10}$ $-5.9112\times 10^{10}$ $1.0590\times 10^{10}$
@@
An NFW dark matter halo (Navarro et al., 1997) is used in the modelling of the Galactic potential:
@@
where $M_{h}=3.3\times 10^{12}$ is the mass of the halo component, $a_{h}=45.02$ kpc is the length scale and $r^{2}=R^{2}+z^{2}$.

### v2

@@ 2.3 COMPAS Implementation of Pulsar Evolution
This study focuses on canonical pulsars, hence we only describe the evolution of spin and magnetic field of pulsars that are not experiencing any mass transfer. Here we follow magnetic braking for an isolated pulsar to spin down, with braking index $n=3$. The pulsar evolution theory implemented in COMPAS is described in Chattopadhyay2020, and is briefly introduced below.
The surface magnetic field strength of isolated pulsars may decay over time. The exact timescale on which this decay occurs (if at all) is uncertain (see further discussion in Section 2.4). The field decay is thought to be due to the Hall effect leading to a redistribution of the magnetic energy, leading to enhanced Ohmic dissipation Romani:1990Natur; Konar:1997MNRAS.284..311K; Konar:1999MNRAS.303..588K; Konar:1999MNRAS.308..795K; PonsGeppert:2007PRL; Aguilera:2008A&A. In COMPAS, for simplicity we assume an exponentially decaying magnetic field of a neutron star similar to many recent studies Kiel:2008MNRAS; Chattopadhyay2020; Dirson2022; ShiNg:2024ApJ. The decay of the magnetic field after it spins down for a time duration of $t$
@@
where $B_{i}$ is the magnetic field strength of the neutron star before it is evolved by time duration $t$ and $B_{f}$ the magnetic field strength afterwards, $B_{\textrm{min}}$ is the minimum magnetic field of a pulsar which is set to be $10^{8}$ G Zhang:2006MNRAS for the entire analysis, and $\tau$ is the time scale of pulsar magnetic field decay.
The rate of change of the spin period of a neutron star over time is known as the spin period derivative, $\dot{P}$ . Assuming spindown due to a rotating magnetic dipole in a vacuum, $\dot{P}$ is given by
@@
where $R$ is the radius of the pulsar (we assume $R=10$ km for all pulsars at all times in this work), $c$ is speed of light, $\alpha$ is the alignment angle between the magnetic and rotational axes, and
$I=(0.237\pm 0.008)MR^{2}[1+4.2\frac{\textrm{M}}{\textrm{M}_{\odot}}\frac{\textrm{km}}{\textrm{R}}+90(\frac{\textrm{M}}{\textrm{M}_{\odot}}\frac{\textrm{km}}{\textrm{R}})^{4}],$ (5)
being the moment of inertia of the pulsar lattimer2005. Substituting $B$ from Eq. 3 into Eq. 4 and integrating over time $t$ gives
$P^{2}=\frac{16\pi^{2}R^{6}\sin^{2}{\alpha}}{3c^{3}I}[B_{\textrm{min}}^{2}t-\tau B_{\textrm{min}}(B_{f}-B_{i})-\frac{\tau}{2}(B^{2}_{f}-B^{2}_{i})]+P_{0}^{2}.$ (6)
$P_{0}$ is the spin period of the pulsar at the beginning of the evolved time duration $t$, and $B_{i}$ and $B_{f}$ are the pulsar surface magnetic field at the beginning and end of the time step as described in Eq. 3. The evolution of $P$ is dependent on $\alpha$ (Eq. 4). As pointed out in other pulsar population synthesis studies (Dirson2022, e.g.,), the evolution of $\alpha$ is also dependent on $B$, as the field decay impacts the evolution of $\alpha$. A simplification in COMPAS is made so that $\alpha$ is set constant and not evolved. This decouples the evolution between PP/$B$ and $\alpha$. Some works have suggested that $\alpha$ tends to decrease as the pulsar ages, and the old pulsar populations have near-alignment between the magnetic field and rotational axes Young:2010MNRAS. However, results from Johnston2020 suggest that large $\alpha$ is needed to model both $\gamma$-ray and radio pulsars well, especially for younger populations. Hence in this work we set $\alpha=\pi/2$. Evolution of alignment will be included in a future version of COMPAS.
@@ 2.4 Free Parameters of COMPAS Runs
In addition to the default binary parameters, we need to propose distributions of pulsar properties at birth for subsequent modelling. Previous pulsar population studies (Faucher-Giguere:2006ApJ; Popov:2012Ap&SS; Szary2014; Johnston2020; Chattopadhyay2020; Dirson2022; Igoshev:2022MNRAS; Du:2024ApJ, e.g.,) have proposed and refined distributions of pulsar properties at birth, such as the surface magnetic field strength and spin period. In this work, we also choose the distributions of these two quantities at birth, as well as magnetic field decay timescale ($\tau_{d}$) to be free parameters.
By default, COMPAS sets the birth distribution of NS magnetic field as a uniform distribution between $10^{10}$ G to $10^{13}$ G, and the birth distribution of NS spin period as a uniform distribution between 10 to 100 ms. These distributions are in general agreement with those observed from young pulsar populations Manchester:2004bp. We exclude extremely high magnetic field strengths with B $10^{15}$ G in the magnetar regime as magnetars are not powered by rotation like typical pulsars, and may have also have a different formation history.
In our default model, we assume that the magnetic field decay time scale $\tau_{d}=500$ Myr Chattopadhyay2021. We choose $\tau_{d}$ as a free parameter since there is no clear consensus on whether the magnetic field of a pulsar decays at all, or how long it takes to decay. For example, semi-analytical studies 1970ApJ...160..979G; 1981JApA....2..315V; 1985MNRAS.213..613L found that the magnetic field decays on the timescale of the order of a million years, with Dirson2022 recently finding $\tau_{d}=4.6\times 10^{5}$ yr. In comparison, other works 1977ApJ...215..885T; 1987A&A...178..143S; Kiel:2008MNRAS have shown the opposite, finding that $\tau_{d}>100$ Myr is needed in order for theoretical models to match the observed catalogue. Recent work incorporating more sophisticated models of magnetic field decay, pulsar beaming and magnetic field alignment find $\tau_{d}$ in the range 1–10 Myr (Graber:2024ApJ; ShiNg:2024ApJ, e.g.,). Given the uncertainty, we let $\tau_{d}$ vary from 1 Myr to 10,000 Myr in some of our models.
@@
We list the details of all the models with different pulsar birth distributions and parameters in Table 1. The initial model (INIT) is as described above. Following that, we vary our model parameters in such ways that we cover different ranges, and in some cases different shapes of the distribution. Models D1–5, combined with INIT, are specifically targeted to examine the impact of the $\tau_{d}$ on the results of the simulations, as only $\tau_{d}$ is varied in models D1–5 compared to INIT. For B1 and B2 models, only the range of the birth magnetic field distributions is changed compared to INIT. Similarly for P1, P2 and P3 models, only the ranges of the birth spin period distributions are changed and the others are kept the same. For PN1, PN2 and PN3 models, the birth spin period distribution is changed to a normal distribution (hence, PN models as in period-normal models) with various means and standard deviations. Model PLN1 gives the birth spin period distribution in a log uniform distribution. In the BLN series of models, the birth magnetic field strength distribution is changed to a log normal distribution, with different means and standard deviations in models BLN1–4. Models P2-D1–5, PN1-B1-D1–5 and PN5-B1-D1–5 are set to explore the impact of $\tau_{d}$ on different initial distributions of spin period and magnetic field, whilst also varying $\tau_{d}$ in the same way as in models D1–5. The remaining three models, BLN1-PN1, BLN4-PN1, and BLN1-PN5, have a combination of the magnetic field distribution from the BL models and the spin period distributions of the PN models, respectively. The original form of the log-normal distribution of the birth magnetic field strength is taken from Szary2014. The BLN1-PN5 model is intended to reproduce the model described in Szary2014. The birth magnetic field and spin period distributions in BLN5-PLN2-D4 are found in Graber:2024ApJ. The birth distributions for BLN6-PLN3-D4 are found in PardoAraujo:2025A&A. In Graber:2024ApJ; PardoAraujo:2025A&A the authors model the magnetic field decay based on numerical simulations of magnetothermal evolution in the neutron star crust Vigano:2021 up to 1 Myr, and with a power law decay model beyond 1 Myr. Their results are equivalent to a short decay timescale. Hence for these two models, the magnetic field decay time scale is set to 1 Myr. In each COMPAS model, 300,000 binary systems are modelled.
@@
Initial ($10^{10}$ - $10^{13}$) Uniform (10 - 100) Uniform 500
D1 ($10^{10}$ - $10^{13}$) Uniform (10 - 100) Uniform 100
D2 ($10^{10}$ - $10^{13}$) Uniform (10 - 100) Uniform 10
D3 ($10^{10}$ - $10^{13}$) Uniform (10 - 100) Uniform 1,000
D4 ($10^{10}$ - $10^{13}$) Uniform (10 - 100) Uniform 1
D5 ($10^{10}$ - $10^{13}$) Uniform (10 - 100) Uniform 10,000
B1 ($10^{11}$ - $10^{13}$) Uniform (10 - 100) Uniform 500
B2 ($10^{10}$ - $10^{12}$) Uniform (10 - 100) Uniform 500
P1 ($10^{10}$ - $10^{13}$) Uniform (10 - 1000) Uniform 500
P2 ($10^{10}$ - $10^{13}$) Uniform (1 - 100) Uniform 500
P3 ($10^{10}$ - $10^{13}$) Uniform (1 - 1000) Uniform 500
PN1 ($10^{10}$ - $10^{13}$) Uniform $\mu=75$$\sigma=25$ Normal 500
PN2 ($10^{10}$ - $10^{13}$) Uniform $\mu=50$$\sigma=25$ Normal 500
PN3 ($10^{10}$ - $10^{13}$) Uniform $\mu=50$$\sigma=50$ Normal 500
PLN1 ($10^{10}$ - $10^{13}$) Uniform $\mu=10^{1.60}$$\sigma=10^{1.40}$ LogNormal 500
@@
P2-D1 ($10^{10}$ - $10^{13}$) Uniform (1 - 100) Uniform 100
P2-D2 ($10^{10}$ - $10^{13}$) Uniform (1 - 100) Uniform 10
P2-D3 ($10^{10}$ - $10^{13}$) Uniform (1 - 100) Uniform 1,000
P2-D4 ($10^{10}$ - $10^{13}$) Uniform (1 - 100) Uniform 1
P2-D5 ($10^{10}$ - $10^{13}$) Uniform (1 - 100) Uniform 10,000
PN1-B1 ($10^{11}$ - $10^{13}$) Uniform $\mu=75$$\sigma=25$ Normal 500
PN1-B1-D1 ($10^{11}$ - $10^{13}$) Uniform $\mu=75$$\sigma=25$ Normal 100
PN1-B1-D2 ($10^{11}$ - $10^{13}$) Uniform $\mu=75$$\sigma=25$ Normal 10
PN1-B1-D3 ($10^{11}$ - $10^{13}$) Uniform $\mu=75$$\sigma=25$ Normal 1,000
PN1-B1-D4 ($10^{11}$ - $10^{13}$) Uniform $\mu=75$$\sigma=25$ Normal 1
PN1-B1-D5 ($10^{11}$ - $10^{13}$) Uniform $\mu=75$$\sigma=25$ Normal 10,000
PN5-B1-D1 ($10^{11}$ - $10^{13}$) Uniform $\mu=300$$\sigma=150$ Normal 100
PN5-B1-D2 ($10^{11}$ - $10^{13}$) Uniform $\mu=300$$\sigma=150$ Normal 10
PN5-B1-D3 ($10^{11}$ - $10^{13}$) Uniform $\mu=300$$\sigma=150$ Normal 1,000
PN5-B1-D4 ($10^{11}$ - $10^{13}$) Uniform $\mu=300$$\sigma=150$ Normal 1
PN5-B1-D5 ($10^{11}$ - $10^{13}$) Uniform $\mu=300$$\sigma=150$ Normal 10,000
BLN1-PN1 $\mu=10^{12.6},\sigma=10^{0.55}$ LogNormal $\mu=75$$\sigma=25$ Normal 500
BLN4-PN1 $\mu=10^{12.0},\sigma=10^{0.25}$ LogNormal $\mu=75$$\sigma=25$ Normal 500
BLN1-PN5 $\mu=10^{12.6},\sigma=10^{0.55}$ LogNormal $\mu=300$$\sigma=150$ Normal 500
BLN5-PLN2-D4 $\mu=10^{13.1},\sigma=10^{0.45}$ LogNormal $\mu=10^{2}$$\sigma=10^{3.38}$ LogNormal 1
BLN6-PLN3-D4 $\mu=10^{13.09},\sigma=10^{0.5}$ LogNormal $\mu=10^{2.33}$$\sigma=10^{3.55}$ LogNormal 1
@@ 2.5 Binary Birth, Pulsar Death, & Parameters
COMPAS begins evolving each binary system at $T=0$ Myr. The first stage of post-processing is to assign a birth time of the binary system that corresponds to the Milky Way star formation history. We consider a uniform star formation history of the Milky Way gomez2018, enabling us to draw the birth time of the binary systems from a uniform distribution. Without recycling, pulsars rapidly spin down, losing their rotational energy to the point that they can no longer sustain detectable radio or $\gamma$-ray emission, at which point we define this as the death of pulsar. The typical characteristic age of a canonical $\gamma$-ray pulsar is relatively small, roughly $10^{5}$ yrs Fermi-LAT:2023zzt. The median characteristic age of radio pulsars selected from the catalogue in this study (as described later in § 3.2) is $3.1\times 10^{6}$ yrs. While this age is higher than typical $\gamma$-ray pulsars, it still suggests that a typical radio pulsar should be formed in a recent star formation episode. For this reason, we assume that all the neutron stars from the simulation that would potentially be $\gamma$-ray emitting are formed in binary systems that are born in the recent history of the Milky Way. Readers should note that in general, the characteristic age is a poor primer for the true age of a pulsar when the neutron star is young or when field decay is included and the star is older than 1 Myr.
The birth time of any given binary system is then drawn from a uniform distribution with a minimum of 12.0 Gyr and a maximum of 13.0 Gyr. We choose to neglect the population of pulsars produced through accretion-induced collapse of white dwarfs as they represent only a small fraction of the population ($\lesssim 10\%$, as indicated by our COMPAS models), are formed on much longer timescales and may potentially be born with different properties to those born in core-collapse supernovae Bailyn:1990ApJ; Tauris:2013A&A.
@@
The emission of a pulsar is turned off when it can no longer produce electron-positron pairs in its magnetosphere. More realistic modelling of pulsar polar cap have suggested a broad “valley” like region in which pulsars cease their emission Chen:1993ApJ; Beskin:2022MNRAS, which is starting to be implemented in some recent pulsar population works (Sautron:2024A&A, e.g.). In this work, we choose a more simplistic description with a death-line model. Before recording the parameters of the neutron star, it is essential to check if the pulsar crosses the radio death-line rudakritter1994 on the $P˙\dot{P}$ diagram, which is defined by
@@
where Eq. 7 is commonly referred to as the death-line for canonical pulsars, and Eq. 8 for MSPs. In the simulation, we decide if a pulsar crosses the death-line (and hence becomes unobservable) by comparing its $\dot{P}$ to the $\dot{P}$ values of Eqs. 7 &  8 given the current spin period of the neutron star. If $\dot{P}$ of a neutron star from COMPAS is larger than both $\dot{P}$ values calculated by Eqs. 7 and  8, then the pulsar has not yet crossed the death-line and may be observable, and its parameters are recorded. Otherwise, it is discarded from the simulation. There are other pulsar population models that do not need to explicitly include a death-line, e.g., when assuming both radio and $\gamma$-ray luminosity are dependent on $\dot{E}$ ShiNg:2024ApJ; Graber:2024ApJ; PardoAraujo:2025A&A. Our treatment of radio selection effects is given in Section 2.7, and $\gamma$-ray selection effects in Section 2.8.
@@ 2.6 Dynamical Evolution
Massive stars, which in later evolutionary stages form neutron stars, are formed within an exponential disk around the Galactic plane. We assume that these massive stars move along their Galactic orbits without any interruption until the change in the evolutionary status of one of the stars in the binary system. As described in § 2.1, we assume all the neutron stars are formed in supernovae. Upon formation, the stellar remnant receives a natal kick as described in § 2.2 which may or may not disrupt the binary system, in some case leading to two individual stars.
Upon the supernova event that forms the neutron star, the binary system is then assigned a random location in the Galaxy following a distribution of a thin, exponential disk Paczynski1990:
$p_{z}(|z|)dz=\exp{(-|z|/z_{\textrm{exp}})}\frac{dz}{z_{\textrm{exp}}},$ (9)
where $|z|$ is the absolute value of the vertical distance from the Galactic plane, and
$p_{R}(R)dR=a_{R}\exp{(-R/R_{\textrm{exp}})}\frac{R}{R^{2}_{\textrm{exp}}}dR,$ (10)
where $a_{R}=1.0683$ and $R$ is the distance from the Galactic centre on the Galactic plane. We set $z_{\textrm{exp}}=75$ pc and $R_{\textrm{exp}}=4.5$ kpc following Paczynski1990 and Dirson2022. Quintana:2025MNRAS recently performed a census of massive OB stars (neutron star progenitors) within 1 kpc of the Sun and found a scale height close to 75 pc. It is worth noting that a recent study of OB stars in Gaia survey Li:2019ApJ has suggested a varying scale height with respect to the Galactic radius. Previous pulsar population synthesis studies have used both larger and smaller scale heights Gullon:2014MNRAS; ShiNg:2024ApJ.
Knowing the location of the newly born neutron star, the time at which it is born, and both the magnitude and direction of the natal kick velocity, we can follow its subsequent motion in the Galactic potential. We use NIGO nigo2015, a numerical galactic orbit integration tool, to evolve the motion of binary systems after the supernova within the Galactic potential. NIGO describes the Milky Way as a disc galaxy, with the Galactic potential comprised of three components: a bulge, a disk and a dark matter halo. The Galactic bulge is described as a Plummer sphere flynn1996; plummer1911:
@@
where $G$ is the gravitation constant, $M_{b}=1.0\times 10^{10}$M is the mass of the bulge and $b_{b}=0.4$kpc is the scale length and $R^{2}=x^{2}+y^{2}$.
The Galactic disc is modelled by an exponential disc formed by the superposition of three Miyamoto-Nagai potentials MN1975:
@@
where $M_{d_{n}}$ are the masses of each disk, $a_{d_{n}}$ are related to the disk scale lengths of the three disk components (which are listed in Table 2) and $b_{d}=0.2376$ is related to the disk scale height.
@@
$M_{d}$ (M) $9.0335\times 10^{10}$ $5.9112\times 10^{10}$ $1.0590\times 10^{10}$
@@
An NFW dark matter halo NFW1997 is used in the modelling of the Galactic potential:
@@
where $M_{h}=3.3\times 10^{12}$M is the mass of the halo component, $a_{h}=45.02$ kpc is the length scale and $r^{2}=R^{2}+z^{2}$.